ICM 1998, Berlin, Aug. 18–27

Abstracts of Short Communications and Posters

Section:

15. Numerical Analysis and Scientific Computing

1991 MS Classification: 65, 73 Wieners, Christian, Universität Stuttgart, Germany Adaptive multigrid methods for Prandtl-Reuß-plasticity We introduce a multigrid method for finite elements and the application to Prandtl-Reuß-plasticity. The basis for the plasticity computations builds a flexible finite element library which supports adaptive multigrid methods for various discretizations. This is coupled by an abstract interface for the material evaluation by a return mapping function at every Gauß-point. We consider in detail the adaptive process of refinement and derefinement and the required interpolation of the material history. The algorithm is realized with the software package UG. This supports parallel multigrid methods on unstructured, locally refined grids with dynamic load balancing. The performance of the algorithm is demonstrated by several examples in 2D with assumed plain strain and in 3D, comparing different material laws (perfect plasticity and exponential hardening).